OpenStudy (anonymous):

Evaluate the expression for the givin value 3q when q=5

OpenStudy (saifoo.khan):

3(5) 15

OpenStudy (anonymous):

f(q)=3q f(3)=3*5 =15

OpenStudy (anonymous):

thats it?

OpenStudy (saifoo.khan):

Oh yes!

OpenStudy (anonymous):

my hero bro

OpenStudy (saifoo.khan):

lol,. ^.^

OpenStudy (bahrom7893):

come on it gave u the q it is asking what is 3 times q.. just multiply whatever q is by 3.. 5*3=15..

OpenStudy (anonymous):

k 8 + w when w = 12 21 right?

OpenStudy (bahrom7893):

8 + 12 = 20 NOT 21

OpenStudy (bahrom7893):

w = 12 so 8+w = 8+12 = 20

OpenStudy (saifoo.khan):

yup!

OpenStudy (anonymous):

kim learning

OpenStudy (anonymous):

take it it easy

OpenStudy (bahrom7893):

oh okay.. but then shouldn't u know that 8+12 is 20...